# Cost-benefit analysis of an AI-driven operational digital platform for integrated electric mobility, renewable energy, and grid management

### v1

@@ 4 Cost-benefit analysis steps with case study
iii) Environmental benefits: Reductions in greenhouse gas (GHG) emissions (primarily CO2) and local air pollutants (SOx, NOx, PM10), contributing to climate change mitigation, improved public health, and greater integration of clean renewable energy.
Each ODP function (A-D) contributes to one or more of these benefit categories. For example, Advanced data collection and processing (A) provides the necessary real-time data for precise grid state estimation and load forecasting, enabling more efficient asset utilization and informed decision-making, which directly translates to economic efficiency and enhanced system reliability. Enhanced interoperability (B) allows for dynamic coordination between different market participants and grid entities, enabling new revenue streams from flexible assets and reducing system imbalances. The core AI-driven forecasting, optimization, and control functions (C) are pivotal in generating significant economic benefits (e.g., reduced energy procurement costs through peak load shaving and optimized energy trading facilitated by highly accurate price predictions, and increased revenues from ancillary service provision). These AI functions also yield substantial reliability benefits (e.g., prevention of grid overloads, improved voltage/frequency stability through predictive control, and reduced RES curtailment by matching demand with variable generation). Furthermore, these AI functions directly contribute to environmental benefits by maximizing the use of RES, optimizing EV charging for lower emission periods, and reducing fossil fuel-based generation for balancing. Decision support and automation (D) translates AI-generated insights into tangible outcomes, yielding direct economic benefits for users (e.g., lower charging costs due to optimized schedules, time savings from efficient routing) and operators (e.g., automated response to grid events), and indirectly fostering environmental benefits by promoting behaviors aligned with RES availability and lower carbon intensity.
@@
Having identified the benefits qualitatively, this step involves quantifying each benefit in monetary terms over the analysis horizon (2026–2035). Specific metrics, mathematical formulations, and assumptions are established for each benefit category. The details are presented in A.1, based upon standard valuation methodologies, relevant literature [11, 40], and country-specific data. The quantification of AI’s contribution is a central focus, linking specific AI model performance metrics and optimization efficiencies to the magnitude of the derived benefits. For each relevant benefit, the improvement margin over the baseline is directly tied to the enhanced capabilities, either superior forecasting accuracy or more efficient operational optimization, provided by the specific AI components of the ODP detailed in Section 3.1. To demonstrate the ODP’s impact in varied national contexts, the proposed CBA framework is applied to Austria (AT), Hungary (HU), and Slovenia (SI). These countries offer a compelling mix of varying grid structures, renewable energy penetration levels, and electric mobility adoption trajectories, providing a representative testbed for the ODP’s cross-border and multi-sectoral capabilities.
@@
| EV stock (2026, thousands) | 343.8 | 100.8 | 50.0 | National Registries, [41] |
| Annual EV Growth rate (CAGR 2026-35) | NECPs, [42] | |||
@@
| ET stock (2026, thousands) | 4.4 | 0.4 | 0.3 | Pilot data, [43] |
@@
| Annual ET growth rate (CAGR 2026-35) | Industry projections, [44] | |||
| RES capacity (Solar + Wind, 2026 GW) | 10.5 | 8.0 | 1.5 | National TSO data, [45] |
| Annual RES addition (MW/yr avg.) | 664 | 2750 | 150 | NECP targets, [46] |
| Avg. electricity retail price (EV, €/kWh) | 0.22 | 0.15 | 0.18 | [47], adj. for EV tariffs |
| Grid CO2 intensity (2026, g/kWh) | 120 | 280 | 190 | [48], projected |
@@
| Value of lost load (VoLL, €/kWh) | 10.0 | 7.5 | 8.0 | Nat. regulator studies, [49] |
@@
| Battery life Ext. (smart charging) | reduction in degradation rate | [37, 38, 36] | ||
@@
| Social cost of carbon (€/tCO2) | Starts at €85 in 2026, rising to €120 by 2035 | EU guidance [50], linear interpolation. | ||
@@
Table 5 summarizes key country-specific input assumptions for the baseline year 2026 and projected growth rates for the analysis period (details are presented in B). These assumptions are derived from recent national energy and climate plans (NECPs), statistical office data [51], reports from international agencies like the IEA [52], and specialized consultancies (e.g., BloombergNEF, McKinsey) to ensure their credibility and relevance to the respective national contexts. For instance, the EV/ET growth rates reflect ambitious but achievable national targets and industry projections, while RES capacity additions align with NECP goals. The assumed reduction in battery degradation rate due to ODP-enabled smart charging, which explicitly incorporates battery health objectives in its AI optimization algorithms (see Section 3.1), is monetized through calculating the net present value of deferred battery replacement costs or the increased residual value of the vehicle. This benefit, derived from the ODP’s AI-driven smart charging (Section 3.1) and validated by internal simulations against established battery aging models [36, 37, 38], involves managing charge/discharge cycles, temperature, and depth of discharge to minimize stress on the battery. Key benefit categories and their monetization approaches are summarized below. Detailed calculations of each benefit and the specific data required for these calculations are presented in Appendices from A.1.1 to A.1.8.
@@
i) Reduction in operational downtime (ROD) cost: This is from avoided power outages or avoided unserved energy due to improved grid reliability and avoided EV/ET service disruptions. A conservative reduction in baseline outage energy could be avoided due to ODP systems enhancement (1) and AI-driven predictive analytics enhancement111For all economic, reliability, and environmental benefits listed here, readers are referred to Section 3.1 for a description of the AI-driven models and the architecture of the ODP that integrates EV, ET, RES, and the grid using different advanced AI layers.
enabling proactive grid management and faster fault restoration. This analysis could be benchmarked against improvements reported in analogous smart grid deployments [53].
@@
iii) Cost savings from demand response and peak load reduction (CSDR-PLR): This is the avoided costs from reducing system peak demand through coordinated EV/ET smart charging and DR programs orchestrated by the ODP’s AI. Peak load reductions of (3) are achieved by the ODP’s MILP-based optimization engine.
@@
v) Fleet energy and operational savings (FES): This is a reduced energy consumption and operational costs for EV/ET fleets. This includes savings from AI-optimized routing (E-VRPTW algorithms in ODP, yielding an estimated reduction in energy/distance for relevant trips, based on pilot study analogs of AI-optimized fleet routing). A significant component is the monetized value of extended battery lifespan (estimated at reduction in degradation rate). This benefit is directly achieved by the ODP’s AI-driven smart charging algorithms, which explicitly incorporate battery health objectives (e.g., optimized SoC profiles, C-rate management, thermal pre-conditioning (6)) informed by established battery aging models [36, 37, 38].
vi) Grid stability and management cost savings (GSMS): This benefit is considered as reduced costs for DSOs/TSOs associated with grid regulation, voltage support, and congestion management. The benefit is calculated as grid stability savings and grid stability from forecasting ability of the AI-driven ODP (7). This estimation is based on the ODP’s AI-driven enhanced forecasting and proactive control capabilities, leading to operational efficiency gains analogous to those reported in advanced smart grid operational studies [53, 54] and analyses of AI-driven grid management platforms.
vii) CO2 emission reduction: This is a monetized benefit of greenhouse gas emission reductions attributable to the ODP’s operations, such as increased RES utilization, optimized EV/ET charging due ODP-enabled RES, EV, ET and awareness creation on the use of operational platform for social cost of carbon reduction (8).
@@
The detailed mathematical formulations and parameter derivations for the above benefits are provided in A.1. By summing the discounted annual benefits across the 10-year analysis period, we obtain the present value for each benefit category and the total benefits.
Applying the monetization formulas with country-specific assumptions yields the projected benefits for each nation.
Table 6 presents the total present value of benefits over , aggregated across the three countries and categorized by type. Detailed annual benefit projections, underpinning these present values, are also provided in B.
@@
Economic benefits constitute the largest share ( of total benefits), primarily driven by the AI-enabled optimized energy trading, demand response, and fleet operational savings. Reliability benefits, quantified mainly through grid stability and management savings (GSMS) at ), presents a significant portion of the benefits, indicating the ODP’s potential to enhance grid resilience and potentially defer traditional grid infrastructure investments. Environmental benefits also show a significant () benefit contribution, showing the ODP’s role in accelerating decarbonization.
Figure 1 illustrates the composition of total benefits, clearly depicting the dominant economic contributions and the significant environmental and reliability gains.
Country-specific analyses (based on the annual data presented in B and Monte Carlo simulation results in A.4) reveal nuanced variations in benefit distribution, reflecting national characteristics. For instance, Austria, with its higher initial EV adoption and ambitious electrification targets, sees proportionally greater benefits from fleet energy and operational savings (FES) and reduction in operational downtime (ROD). Hungary, with its significant RES expansion plans, benefits more substantially from avoided energy curtailment (AEC) and revenue from optimized energy trading and ancillary services (ROETAS) due to the greater flexibility needs. Slovenia, given its smaller scale but gains particularly from optimized ET logistics and enhanced cross-border grid balancing, which the ODP is designed to facilitate. These variations underscore the ODP’s adaptable value proposition across diverse EU member states.
@@
This step involves a comprehensive enumeration and quantification of all costs associated with the implementation and operation of the ODP over the 2026–2035 period. Costs are categorized into Capital expenditures (CAPEX) and Operational expenditures (OPEX). Detailed breakdowns, differentiating between initial platform development and ongoing integration costs, are provided in A.2 (CAPEX) and A.3 (OPEX). Below, we explain their main elements:
CAPEX: These are primarily upfront investments incurred during the initial deployment phase (2026) and for system expansion in subsequent years. Major CAPEX components include hardware procurement (sensors, servers, communication devices), core software development for the ODP platform, and crucially, the initial development and training of the AI engine’s specialized models (e.g., LSTM architectures, MILP/RL solver integration, data pipelines for AI). System integration and project management costs are also included. A significant portion of CAPEX is incremental, related to the integration of new EVs, ETs, and RES installations annually into the ODP-managed ecosystem. Our estimates indicate a total discounted CAPEX of €561.3 million over the 10-year period, with approximately of the initial platform CAPEX directly attributable to AI model development and associated infrastructure (e.g., specialized GPU clusters for training).
OPEX: These are recurring annual costs necessary to operate and maintain the ODP. Key OPEX categories include cloud computing resources for AI inference and data processing, ongoing software maintenance (including continuous AI model monitoring, re-training pipelines to counter data/concept drift, and performance optimization), communication fees, robust cybersecurity operations, labor costs for technical staff (e.g., data scientists, AI engineers, system administrators), and hardware maintenance. Annual OPEX is projected to start at approximately €31.6 million (discounted) in the first year (2026), with an estimated of this directly related to AI operations and maintenance. The total discounted OPEX over 10 years is estimated at €315.9 million.
@@
The total discounted investment cost (CAPEX + OPEX) for the ODP deployment and operation over the decade is approximately €877.2 million. Cost estimates are benchmarked against similar smart grid and digital platform deployments in the energy sector [12, 55] and include reasonable contingencies for unforeseen expenditures. The specific breakdown for AI-related costs captures both initial development (CAPEX) and ongoing operational and maintenance expenses (OPEX) essential for sustaining AI model performance in a dynamic environment.
Total costs (CAPEX and OPEX) for the aggregated three-country deployment amount to €877.2 million. CAPEX for shared platform development and core infrastructure, including AI engine development, was allocated to each country based on a composite weighting factor considering projected deployment scale (e.g., number of managed EVs/ETs, RES capacity integrated, and overall energy consumption), as detailed in A.2 and A.3. As a result, Austria accounts for approximately of total costs, Hungary , and Slovenia . Detailed cost breakdowns, including specific required items for AI system development and ongoing AI model maintenance, are comprehensively presented in A.2 and A.3, with annual cost projections provided in B.
@@
In this step, the total present value of monetized benefits is compared against the total present value of costs. Key financial performance indicators, including NPV and BCR, are calculated to assess the overall economic viability of the ODP investment. Table 7 summarizes the aggregated financial outcomes over the 10-year analysis horizon (2026–2035), discounted at .
@@
The positive NPV and BCR 1 strongly support the economic and financial feasibility for the ODP deployment. These results are based on conservative assumptions for benefit realization and cost estimations, detailed in A.2 and A.3, and include the explicit quantification of AI’s performance-driven contributions.
@@
| Total benefits + | ||
| Total benefits - | ||
| Total costs + | ||
| Total costs - | ||
| Discount rate (instead of ) | ||
| Discount rate (instead of ) |
@@
The scenario analysis indicates that the ODP investment remains viable (BCR 1, NPV 0) even under reasonably pessimistic assumptions. The sensitivity analysis demonstrates a comfortable margin of safety for the investment. A Monte Carlo simulation ( iterations) was also performed. Key parameters such as EV/ET adoption rates, energy price volatility, RES generation variability, CAPEX/OPEX unit costs, AI model performance factors (e.g., ranges for forecast accuracy benefits, optimization efficiency gains), and the social cost of carbon were modeled as random variables with appropriate probability distributions (e.g., triangular or normal distributions based on expert judgment and historical variability). The specific ranges and distributions were derived from a range of possible real-world outcomes (e.g., slightly lower forecast accuracy than optimal, or less efficient optimization in unforeseen scenarios).
@@
Figure 4 presents the resulting probability distributions for NPV and BCR, providing a probabilistic view of the project’s financial outcomes under uncertainty.
@@
The Monte Carlo simulation indicates a mean NPV of approximately €365 million and a mean BCR of 1.41. The probability of achieving an NPV greater than zero is estimated to be over , while the probability of a BCR greater than 1.0 is also over . The slight difference between the Monte Carlo mean NPV/BCR and the deterministic base case values can arise from asymmetric probability distributions of input variables, the presence of significant non-linear relationships between input variables and the output metrics, or correlations between input variables, which are inherently captured in a probabilistic approach. These results strongly reinforce the financial robustness of the ODP investment under uncertainty, demonstrating high confidence in its viability even when considering potential variations in adoptions and discount rates, and other key parameters.
@@
Qualitative risk assessment identified additional factors: regulatory and policy risks (e.g., changes in market design or data sharing regulations), cybersecurity and data privacy risks (given the sensitive nature of energy and mobility data), technology adoption and integration risks (e.g., slower than expected stakeholder uptake, challenges in integrating with legacy systems), and stakeholder acceptance. Mitigation strategies include proactive engagement with policymakers to shape supportive regulatory frameworks, robust cybersecurity measures adhering to international standards like ISO 27001 and GDPR compliance, a phased deployment approach to manage technical complexities, and a user-centric design philosophy to enhance stakeholder acceptance.
The sensitivity patterns observed in the aggregated analysis (Section 4) generally hold true at the country level. This indicates that while the absolute values of benefits and costs vary, the project’s overall viability and robustness to uncertainties remain consistent across different national contexts. Monte Carlo simulations performed independently for each country consistently indicate a probability of achieving a BCR 1 exceeding in all three national cases. This further reinforces the economic rationale for the ODP’s deployment at both regional and national scales, providing a high degree of confidence in its financial success despite inherent market and technological uncertainties.

### v2

@@ 4 Cost-benefit analysis steps with case study
iii) Environmental benefits: Reductions in greenhouse gas (GHG) emissions (primarily CO2) and local air pollutants (SOx, NOx, PM2.5), contributing to climate change mitigation, improved public health, and greater integration of clean renewable energy.
Each ODP function (A-D) contributes to one or more of these benefit categories. For example, advanced data collection and processing (A) provides the necessary real-time data for precise grid state estimation and load forecasting, enabling more efficient asset utilization and informed decision-making, which directly translates to economic efficiency and enhanced system reliability. Enhanced interoperability (B) enables dynamic coordination between different market participants and grid entities, enabling new revenue streams from flexible assets and reducing system imbalances. The core AI-driven forecasting, optimization, and control functions (C) are pivotal in generating significant economic benefits (e.g., reduced energy procurement costs through peak load shaving and optimized energy trading facilitated by highly accurate price predictions, and increased revenues from ancillary service provision). These AI functions also yield substantial reliability benefits (e.g., prevention of grid overloads, improved voltage/frequency stability through predictive control, and reduced RES curtailment by matching demand with variable generation). Furthermore, these AI functions directly contribute to environmental benefits by maximizing the use of RES, optimizing EV charging for lower emission periods, and reducing fossil fuel-based generation for balancing. Decision support and automation (D) translates AI-generated insights into tangible outcomes, yielding direct economic benefits for users (e.g., lower charging costs due to optimized schedules, time savings from efficient routing) and operators (e.g., automated response to grid events), and indirectly fostering environmental benefits by promoting behaviors aligned with RES availability and lower carbon intensity.
@@
As described in Step 1, all benefits of the ODP-enhanced system are measured as the difference between the system with ODP and a counterfactual baseline (without ODP). This is done by defining a physical delta relative to the baseline and applying a valuation.
where is a monetary valuation of the delta-benefit, such as price.
The CBA adopts a 10‐year appraisal window (2026–2035) appropriate for a AI‐centric ODP whose enabling technologies undergo major upgrades within 5–10 years. This aligns with prior digital/smart‐grid CBAs Vilaplana et al. [2025], Ruutu et al. [2017] and avoids speculative long‐term assumptions typical of structural assets. A 15‐year sensitivity (2026–2040) using the same trajectories—with a mid‐life refresh—confirms the robustness of conclusions and yields higher NPV and BCR. Specific metrics, mathematical formulations, and assumptions are established for each benefit category. The details are presented in A.1, based upon standard valuation methodologies, relevant literature [EU Commission JRC, 2012, BloombergNEF, 2022], and country-specific data. For each relevant benefit, the improvement margin over the baseline is directly tied to the enhanced capabilities, either superior forecasting accuracy or more efficient operational optimization, provided by the specific AI components of the ODP. To demonstrate the ODP’s impact in varied national contexts, the proposed CBA framework is applied to Austria (AT), Hungary (HU), and Slovenia (SI). These countries offer a compelling mix of varying grid structures, renewable energy penetration levels, and electric mobility adoption trajectories, providing a representative testbed for the ODP’s cross-border and multi-sectoral capabilities.
@@
| EV stock (2026, thousands) | 343.8 | 100.8 | 50.0 | National Registries, [Statista Research Department, 2023] |
| Annual EV Growth rate (CAGR 2026-35) | NECPs, [BloombergNEF, 2023] | |||
@@
| ET stock (2026, thousands) | 4.4 | 0.4 | 0.3 | Pilot data, [European Automobile Manufacturers Association (2023), ACEA] |
@@
| Annual ET growth rate (CAGR 2026-35) | Industry projections, [McKinsey & Company, 2022] | |||
| RES capacity (Solar + Wind, 2026 GW) | 10.5 | 8.0 | 1.5 | National TSO data, [ENTSO-E, 2022] |
| Annual RES addition (MW/yr avg.) | 664 | 2750 | 150 | NECP targets, [SolarPower Europe, 2023] |
| Avg. electricity retail price (EV, €/kWh) | 0.22 | 0.15 | 0.18 | [Eurostat, 2023], adj. for EV tariffs |
| Grid CO2 intensity (2026, g/kWh) | 120 | 280 | 190 | [European Environment Agency (2023), EEA], projected |
@@
| Value of lost load (VoLL, €/kWh) | 10.0 | 7.5 | 8.0 | Nat. regulator studies, [Agency for the Cooperation of Energy Regulators (2021), ACER] |
@@
| Battery life Ext. (smart charging) | reduction in degradation rate | [Curry et al., 2023, Maheshwari et al., 2020] | ||
@@
| Social cost of carbon (€/tCO2) | Starts at €85 in 2026, rising to €120 by 2035 | EU guidance [European Commission, 2022], linear interpolation. | ||
@@
Table 5 summarizes key country-specific input assumptions for the baseline year 2026 and projected growth rates for the analysis period (details are presented in B). These assumptions are derived from recent national energy and climate plans (NECPs), statistical office data [Eurostat, 2023], reports from international agencies like the IEA [International Energy Agency (2023), IEA], and specialized consultancies (e.g., BloombergNEF, McKinsey) to ensure their credibility and relevance to the respective national contexts. For instance, the EV/ET growth rates reflect ambitious but achievable national targets and industry projections, while RES capacity additions align with NECP goals. Battery-health benefits are quantified using an explicit aging representation within the ODP architecture and monetized as the NPV of the modeled reduction in degradation cost relative to a baseline without ODP. Empirical studies report an 8-12% lower degradation under health-aware charging, narrower state-of-charge windows, C-rate limits, temperature-aware scheduling, and reduced high-SoC dwell, compared with uncoordinated practice [Maheshwari et al., 2020, Curry et al., 2023]. In this study, the empirical range serves as an external validation band lower bound (8%), which is used in the calculation as a conservative figure.
Key benefit categories and their monetization approaches are summarized below. Detailed calculations of each benefit and the specific data required for these calculations are presented in Appendices from A.1.1 to A.1.8.
@@
i) Reduction in operational downtime (ROD) cost: This is from avoided power outages or avoided unserved energy due to improved grid reliability and avoided EV/ET service disruptions. A conservative reduction in baseline outage energy could be avoided due to ODP systems enhancement (1) and AI-driven predictive analytics enables proactive grid management and faster fault restoration. This analysis could be benchmarked against improvements reported in analogous smart grid deployments [U.S. Department of Energy, Office of
Electricity, 2019].
@@
iii) Cost savings from demand response and peak load reduction (CSDR-PLR): This is the avoided costs from reducing system peak demand through coordinated EV/ET smart charging and DR programs orchestrated by the ODP’s AI. Peak load reductions of (3) are achieved by the ODP’s optimization engine.
@@
v) Fleet energy and operational savings (FES): This is a reduced energy consumption and operational costs for EV/ET fleets. This includes savings from AI-optimized routing (E-VRPTW algorithms in ODP, yielding an estimated reduction in energy/distance for relevant trips, based on pilot study analogs of AI-optimized fleet routing). A significant component is the monetized value of extended battery lifespan, which is estimated from 8–12% reduction in degradation rate Maheshwari et al. [2020].
This benefit is directly achieved by the ODP’s AI-driven smart charging algorithms, which explicitly incorporate battery health objectives informed by established battery aging models [Maheshwari et al., 2020, Curry et al., 2023].
vi) Grid stability and management cost savings (GSMS): This benefit is considered as reduced costs for DSOs/TSOs associated with grid regulation, voltage support, and congestion management. The benefit is calculated as grid stability savings and grid stability from the forecasting ability of the AI-driven ODP (7). This estimation is based on the ODP’s AI-driven enhanced forecasting and proactive control capabilities, leading to operational efficiency gains analogous to those reported in advanced smart grid operational studies [U.S. Department of Energy, Office of
Electricity, 2019, Electric Power Research Institute
(2020), EPRI] and analyses of AI-driven grid management platforms.
vii) CO2 emission reduction: This is a monetized benefit of greenhouse gas emission reductions attributable to the ODP’s operations, such as increased RES utilization, optimized EV/ET charging due ODP-enabled RES, EV, ET and awareness creation on the use of the operational platform for social cost of carbon reduction.
@@
The detailed mathematical formulations and parameter derivations for the above benefits are provided in A.1. By summing the discounted annual benefits across the 10-year period, we calculate the present value for each benefit category and the total benefits. Table 6 presents the total present value of benefits over , aggregated across the three countries and categorized by type. Detailed annual benefit projections, underpinning these present values, are also provided in B.
@@
Economic benefits constitute the largest share ( of total benefits), primarily driven by the AI-enabled optimized energy trading, demand response, and fleet operational savings. Reliability benefits, quantified mainly through grid stability and management savings (GSMS) at ), presents a significant portion of the benefits, indicating the ODP’s potential to enhance grid resilience and potentially defer traditional grid infrastructure investments. Environmental benefits also contribute significantly () highlighting the ODP’s role in accelerating decarbonization.
This enhancement increases the granularity and policy relevance of the environmental assessment without altering the fundamental framework for benefit calculation. Figure 3 presents total benefits by category and their year-by-year evolution. The top-left panel displays the total economic, reliability, and environmental benefits; the top-right panel shows the annual trajectory of each benefit over the project horizon. The bottom panel plots the air-pollutant trends under the existing system and the ODP-enhanced system.
The right panel depicts the yearly trajectory of each benefit in the project horizon. Specifically, country-specific analyses (based on the annual data presented in B and Monte Carlo simulation results in Table 4) reveal nuanced variations in benefit distribution, reflecting national characteristics. For instance, Austria, with its higher initial EV adoption and ambitious electrification targets, sees proportionally greater benefits from fleet energy savings (FES) and a reduction in operational downtime (ROD). Hungary, with its significant RES expansion plans, benefits more substantially from avoided energy curtailment (AEC) and revenue from optimized energy trading and ancillary services (ROETAS) due to the greater flexibility needs. Slovenia, given its smaller scale but gains particularly from optimized ET logistics and enhanced cross-border grid balancing, which the ODP is designed to facilitate. These variations underscore the ODP’s adaptable value proposition across diverse EU member states.
@@
This step involves a comprehensive enumeration and quantification of all costs associated with implementing and operating the ODP over the 2026–2035 period. Costs are categorized into Capital expenditures (CAPEX) and Operational expenditures (OPEX). Detailed breakdowns, differentiating between initial platform development and ongoing integration costs, which are provided in A.2 (CAPEX) and A.3 (OPEX). Below, we explain their main elements:
CAPEX: These are primarily upfront investments incurred during the initial deployment phase (2026) and for system expansion in subsequent years. Major CAPEX components include hardware procurement (sensors, servers, communication devices), core software development for the ODP platform, and crucially, the initial development and training of the AI engine’s specialized models. System integration and project management costs are also included. A significant portion of CAPEX is incremental, related to the integration of new EVs, ETs, and RES installations annually into the ODP-managed ecosystem. Our estimates indicate a total discounted CAPEX of €561.3 million over the 10-year period, with approximately of the initial platform CAPEX directly attributable to AI model development and associated infrastructure (e.g., specialized GPU clusters for training).
OPEX: These are recurring annual costs necessary to operate and maintain the ODP. Key OPEX categories include cloud computing resources for AI inference and data processing, ongoing software maintenance (including continuous AI model monitoring, re-training pipelines to counter data/concept drift, and performance optimization), communication fees, robust cybersecurity operations, labor costs for technical staff (e.g., data scientists, AI engineers, system administrators), and hardware maintenance. Annual OPEX is projected to start at approximately €31.6 million (discounted) in the first year (2026), with an estimated 20% of this directly related to AI operations and maintenance. The total discounted OPEX over 10 years is estimated at €315.9 million.
@@
The total discounted investment cost (CAPEX + OPEX) for the ODP deployment and operation over the decade is approximately €877.2 million. Cost estimates are benchmarked against similar smart grid and digital platform deployments in the energy sector [Stromsather et al., 2018, ETIP SNET, 2022] and include reasonable contingencies for unforeseen expenditures. The specific breakdown for AI-related costs captures both initial development (CAPEX) and ongoing operational and maintenance expenses (OPEX), which are essential for sustaining the performance of AI models in a dynamic environment.
CAPEX for shared platform development and core infrastructure, including AI engine development, was allocated to each country based on a composite weighting factor considering projected deployment scale (e.g., number of managed EVs/ETs, RES capacity integrated, and overall energy consumption), as detailed in A.2 and A.3. As a result, Austria accounts for approximately 45% of total costs, Hungary 35%, and Slovenia 20%. Detailed cost breakdowns, including specific required items for AI system development and ongoing AI model maintenance, are comprehensively presented in A.2 and A.3, with annual cost projections provided in B.
@@
In this step, the total present value of monetized benefits is compared against the total present value of costs. Key financial performance indicators, including NPV and BCR, are calculated to assess the overall economic viability of the ODP investment. Table 7 summarizes the aggregated financial outcomes over the 10-year analysis horizon, discounted at 4%.
@@
The positive NPV and BCR 1 strongly support the economic and financial feasibility of the ODP deployment. These results are based on conservative assumptions for benefit realization and cost estimations, detailed in A.2 and A.3, and include the explicit quantification of AI’s performance-driven contributions.
@@
| Total benefits +10% | ||
| Total benefits -10% | ||
| Total costs +10% | ||
| Total costs -10% | ||
| Discount rate 3% (instead of 4% | ||
| Discount rate 5% (instead of 4%) |
@@
The scenario analysis indicates that the ODP investment remains viable (BCR 1, NPV 0) even under reasonably pessimistic assumptions. The sensitivity analysis demonstrates a comfortable margin of safety for the investment. A Monte Carlo simulation ( iterations) was also performed as a probabilistic risk assessment framework. Key parameters such as EV/ET adoption rates, energy price volatility, CAPEX/OPEX unit costs, AI model performance factors (e.g., ranges for forecast accuracy benefits, optimization efficiency gains), and data availability are considered. The specific ranges and distributions were derived from a range of possible real-world outcomes (e.g., slightly lower forecast accuracy than optimal, or less efficient optimization in unforeseen scenarios).
@@
Figure 6 shows the probability distributions of NPV and BCR, providing a probabilistic view of the project’s financial outcomes under uncertainty. The simulations vary energy-price volatility, EV/ET adoption rates, and CAPEX/OPEX around the baseline NPV and BCR estimates.
@@
The Monte Carlo simulation indicates a mean NPV of approximately €365 million and a mean BCR of 1.41. The probability of achieving an NPV greater than zero is estimated to be over , while the probability of a BCR greater than 1.0 is also over . The slight difference between the Monte Carlo mean NPV/BCR and the deterministic base case values can arise from asymmetric probability distributions of input variables, the presence of significant non-linear relationships between input variables and the output metrics, or correlations between input variables, which are inherently captured in a probabilistic approach. These results strongly reinforce the financial robustness of the ODP investment under uncertainty, demonstrating high confidence in its viability even when considering potential variations in adoption and discount rates, as well as other key parameters.
Figure 7 presents the extended sensitivity analysis focusing on four critical parameters: CAPEX, OPEX, data availability, and AI accuracy. The results indicate that NPV ranges from approximately € 100 million to € 700 million, while BCR varies between 1.0 and 1.8. CAPEX assumptions exert the strongest influence on NPV, followed by AI accuracy and data availability, both of which play a crucial role in determining the economic viability of the ODP. These underscore that, although the framework is sensitive to technological performance and data quality, the ODP consistently shows positive value across the explored scenarios.
@@
Qualitative risk assessment identified additional factors: regulatory and policy risks (e.g., changes in market design or data sharing regulations), cybersecurity and data privacy risks (given the sensitive nature of energy and mobility data), technology adoption and integration risks (e.g., slower than expected stakeholder uptake, challenges in integrating with legacy systems), and stakeholder acceptance. Mitigation strategies include proactive engagement with policymakers to shape supportive regulatory frameworks, robust cybersecurity measures that adhere to international standards such as ISO 27001 and GDPR compliance, a phased deployment approach to manage technical complexities, and a user-centric design philosophy to enhance stakeholder acceptance. The sensitivity patterns observed in the aggregated analysis generally hold true at the country level. This suggests that although the absolute values of benefits and costs vary, the project’s overall viability and robustness to uncertainties remain consistent across different national contexts. Monte Carlo simulations performed independently for each country consistently indicate a probability of achieving a BCR 1 exceeding in all three national cases. This further reinforces the economic rationale for the ODP’s deployment at both regional and national scales, providing a high degree of confidence in its financial success despite inherent market and technological uncertainties.
Per-country CBA summary and Monte Carlo simulation (MCS) results
Now, we can see a summary of the CBA outcomes and Monte Carlo simulation (MCS) results for the AI-driven ODP, disaggregated for each of the three countries: Austria (AT), Hungary (HU), and Slovenia (SI). The numerical values presented in Table 9 represent the actual calculated results, based on country-specific assumptions and scaled costs, which demonstrate the national-level impacts derived from the detailed financial and operational models.
| Metric | Austria (AT) | Hungary (HU) | Slovenia (SI) |
| Deterministic results (Base Case) | |||
| Total benefits (PV) | 550.8 | 434.7 | 248.4 |
| Total costs (PV) | 390.4 | 310.8 | 176.0 |
| NPV (PV) | 160.4 | 123.9 | 72.4 |
| BCR | 1.41 | 1.40 | 1.41 |
| Monte Carlo simulation (Mean values from 10,000 Trials) | |||
| Mean NPV (PV) | 168.2 | 127.5 | 74.3 |
| Mean BCR | 1.43 | 1.42 | 1.42 |
| Dominant benefit categories (% of national total benefits) | |||
| Economic (ROD, AEC, FES, ROETAS, CSDR-PLR) | |||
| Environmental (CO2, RAP) | |||
| Reliability (GSMS) | |||
1.
Note: These numerical values are derived from detailed country-specific financial and operational models, reflecting the specific growth rates, market prices, and grid characteristics outlined in Table 5 and other appendices. The slight variations in BCR across countries reflect the unique cost structures and benefit profiles of each national context. Dominant benefit categories highlight the primary drivers of value in each country, influenced by the respective characteristics of their energy and transport systems.